0.10 m soultions of sodium hydroxide

7.48 Part A What is the pH of a solution that has a hydroxide ion concentration of 1.0 × 10-7 M? pH = Type here to searc
Answer #1

Here, [OH-] = 10-7 M

pOH = -Log[OH-]

= -Log(10-7)

= 7

pH = 14 - pOH

= 14 - 7

= 7

Therefore, the pH of a solution that has a hydroxide ion concentration of 1.0*10-7 M = 7

Part (ii): The pOH of 0.1 M solution of NaOH = -Log(0.1) = 1

i.e. pH = 14 - 1 = 13